### v1

# Uniform rank metric stability of Lie algebras, Lie groups and lattices

Abstract.
We study uniform stability of discrete groups, Lie groups and Lie algebras in the rank metric, and the connections between uniform stability of these objects. We prove that semisimple Lie algebras are far from being flexibly $\mathbb{C}$-stable, and that semisimple Lie groups and lattices in semisimple Lie groups of higher rank are not strictly $\mathbb{C}$-stable. Furthermore, we prove that free groups are not uniformly flexibly $F$-stable over any field $F$.
1. Introduction
The notion of homomorphism stability is the general problem of whether almost homomorphisms are close to genuine homomorphisms, and is usually referred to as Ulam stability [16]. This property depends on the objects in question (usually groups or algebras), and on the precise meaning of almost homomorphisms and distances between them, as well as the chosen metric. A theorem of Kazhdan gives such a result for unitary almost representations in the operator norm:
Theorem.
(Kazhdan 1982 [11])
Let $\Gamma$ be an amenable group and $\varphi:\Gamma\to U\left(\mathcal{H}\right)$ be a function into the unitary operators on a Hilbert space $\mathcal{H}$. If $\left\|\varphi\left(\gamma_{1}\gamma_{2}\right)-\varphi\left(\gamma_{1}\right)\varphi\left(\gamma_{2}\right)\right\|_{op}\leq\delta$ for all $\gamma_{1},\gamma_{2}\in\Gamma$ and $\delta\leq\frac{1}{200}$, then there exists a representation $\psi:\Gamma\to U\left(\mathcal{H}\right)$ such that $\left\|\varphi\left(\gamma\right)-\psi\left(\gamma\right)\right\|_{op}\leq 2\delta$ for all $\gamma\in\Gamma$.
In general, we say that the group $\Gamma$ is uniformly stable in the operator norm, if for any $\varepsilon>0$ there exists a $\delta>0$ such that for any $n\in\mathbb{N}$ and any map $\varphi:\Gamma\to U\left(n\right)$ satisfying
$\left\|\varphi\left(\gamma_{1}\gamma_{2}\right)-\varphi\left(\gamma_{1}\right)\varphi\left(\gamma_{2}\right)\right\|_{op}\leq\delta\quad\forall\gamma_{1},\gamma_{2}\in\Gamma,$
there exists a genuine representation $\psi:\Gamma\to U\left(n\right)$ such that $\left\|\varphi\left(\gamma\right)-\psi\left(\gamma\right)\right\|_{op}\leq\varepsilon$ for all $\gamma\in\Gamma$ [^1] This property was called “strong Ulam stability” in [5]..
Burger, Ozawa and Thom [5] have proved that the groups $SL_{n}\left(\mathcal{O}_{S}\right)$ are uniformly stable in the operator norm, where $n\geq 3$ and $\mathcal{O}$ is the ring of integers of a number field and $\mathcal{O}_{S}$ some localization of $\mathcal{O}$. Glebsky, Lubotzky, Monod and Rangarajan [8] have proved uniform stability in $U\left(n\right)$ with respect to arbitrary submultiplicative norms for all amenable groups (generalizing Kazhdan’s theorem for the case of finite dimensional Hilbert spaces) and for a large class of lattices in higher rank semisimple Lie groups using vanishing properties of asymptotic cohomology. Rolli [15] proved a negative result for nonabelian free groups by constructing maps $\varphi:\mathbb{F}_{m}\to U\left(n\right)$ with $\left\|\varphi\left(\gamma_{1}\gamma_{2}\right)-\varphi\left(\gamma_{1}\right)\varphi\left(\gamma_{2}\right)\right\|_{op}\leq\delta$ for any $\gamma_{1},\gamma_{2}\in\Gamma$ and arbitrarily small $\delta$, that are bounded from all representations $\psi:\Gamma\to U\left(n\right)$.
Following Arzhantseva and Păunescu [2], we consider a similar problem with the unitary groups replaced by $GL_{n}\left(F\right)$ (for some field $F$), endowed with the normalized rank metric:
$rk\left(A-B\right)\vcentcolon=\frac{1}{n}rank\left(A-B\right)\qquad\forall A,B\in GL_{n}\left(F\right).$
This metric can be seen as a generalization of the normalized Hamming metric from finite permutation groups to general linear groups, see [2] for details.
For a group $\Gamma$, an $\varepsilon$-representation is a function $\varphi:\Gamma\to GL_{n}\left(F\right)$ satisfying
[…]
We say that the group $\Gamma$ is uniformly $F$-stable if for any $\varepsilon>0$, there exists a $\delta>0$ such that for any $n\in\mathbb{N}$ and any $\delta$-representation $\varphi:\Gamma\to GL_{n}\left(F\right)$, there exists a representation $\psi:\Gamma\to GL_{n}\left(F\right)$ such that $rk\left(\varphi\left(\gamma\right)-\psi\left(\gamma\right)\right)\leq\varepsilon\quad\forall\gamma\in\Gamma$.
Our first main result shows that many lattices in higher rank semisimple Lie groups are not uniformly $\mathbb{C}$-stable, as opposed to submultiplicative norms as in [8].
Theorem 1.1.
Let $G$ be a connected linear semisimple real Lie group, defined over $\mathbb{Q}$ and of rank at least $2$ with no compact factors, and $\Gamma\leq G$ an irreducible lattice. Then $\Gamma$ is not uniformly $\mathbb{C}$-stable.
Similarly to stability in permutations [4] and in the normalized Hilbert-Schmidt norm [6], we consider also a notion of flexible stability, where the homomorphism $\psi:\Gamma\to GL_{N}\left(F\right)$ is allowed to take values in matrices of larger dimension $N\geq n$, and $\varphi\left(\Gamma\right)$ is embedded in $M_{N}\left(F\right)$ as a corner, see Section 2.1. We will usually refer to stability in the normal sense (where an increase in dimension is not allowed) as strict stability.
The question of whether lattices as in Theorem 1.1 are flexibly uniformly stable remains open, and we discuss this in Section 3.3.
Our second main result concerns flexible stability of semisimple Lie algebras, which will be shown to be connected to the stability of semisimple Lie groups and their lattices. For a Lie algebra $\mathfrak{g}$, an $\varepsilon$-representation is a linear map $\varphi:\mathfrak{g}\to\mathfrak{gl}_{n}\left(F\right)$ satisfying
$rk\left(\left[\varphi\left(x\right),\varphi\left(y\right)\right]-\varphi\left(\left[x,y\right]\right)\right)\leq\varepsilon\quad\forall x,y\in\mathfrak{g},$
and $\mathfrak{g}$ is uniformly $F$-stable if for any $\varepsilon>0$, there exists a $\delta>0$ such that for any $n\in\mathbb{N}$ and any $\delta$-representation $\varphi:\mathfrak{g}\to\mathfrak{gl}_{n}\left(F\right)$, there exists a representation $\psi:\mathfrak{g}\to\mathfrak{gl}_{n}\left(F\right)$ such that $rk\left(\varphi\left(x\right)-\psi\left(x\right)\right)\leq\varepsilon\quad\forall x\in\mathfrak{g}$.
The following result was proven independently in [3].
Theorem 1.2.
Let $\mathfrak{g}$ be a finite-dimensional semisimple real Lie algebra. Then $\mathfrak{g}$ is not flexibly uniformly $\mathbb{C}$-stable.
Finally, we provide counterexamples for flexible stablity of groups, by a construction in the spirit of Rolli [15] who proved that free groups are not stable in the operator norm, in addition to the proof in [4] that free groups are not flexibly stable in permutations.
Theorem 1.3.
Let $\Gamma$ be a group that surjects onto the free group $\mathbb{F}_{2}$ of rank $2$. Then $\Gamma$ is not flexibly $F$-stable for any field $F$.
There are various results on uniform stability of discrete groups with respect to different classes of groups and metrics. De-Chiffre, Ozawa and Thom [6] have generalized a result by Gowers and Hatami [9], proving that all amenable groups are uniformly flexibly stable in the Hilbert-Schmidt norm. In contrast, Akhtiamov and Dogon [1] have proved that for $\Gamma$ a finitely generated residually finite group, strict uniform Hilbert-Schmidt stability is equivalent to $\Gamma$ being virtually abelian.
Becker and Chapman [4] have proved that amenable groups are uniformly flexibly stable in permutations, and that in many cases the flexibility is necessary; For any group $\Gamma$ with finite quotients of unbounded size, $\Gamma$ is not strictly stable. Additionally, they have proved that the groups $SL_{r}\left(\mathcal{O}_{K}\right)$ are uniformly flexibly stable in permutations, where $r\geq 3$ and $\mathcal{O}_{K}$ the ring of integers in a number field $K$.
In general, representations of a group into $GL_{n}\left(\mathbb{C}\right)$ tend to be much less understood than unitary representations. For this reason, we focus in this paper on a class of groups whose (finite dimensional) representation theory is relatively well behaved, namely, irreducible lattices in higher rank semisimple Lie groups. We consider compressions of representations of $\Gamma\leq G$ coming from the Lie group $G$, and taking differentials at the identity we obtain almost representations of its Lie algebra $\mathfrak{g}$. Thus, the notion of stability of Lie algebras arises naturally in this context, and our main focus will be in the case where the Lie group/algebra is semisimple. In Theorem 3.1, we use central characters of representations of $\mathfrak{g}$ to prove the existence of an uncountable family of “highest weight” $\varepsilon$-representations of any semisimple Lie algebra, that are nonequivalent in a suitable sense. In particular, this reestablishes the fact that they are not flexibly stable. We then use the connection between stability of Lie groups and stability of their Lie algebras, as well as rigidity properties for lattices in semisimple Lie groups to deduce Theorem 1.1.
There is an important related notion of pointwise stability, where for a finitely presented group/algebra, an $\varepsilon$-homomorphism is a map $\varphi:S\to GL_{n}\left(F\right)$ ($\mathfrak{gl}_{n}\left(F\right)$ in the case of algebras) from a set of generators $S$ such that the finite set of relations $R$ are $\varepsilon$-satisfied in $GL_{n}\left(F\right)$ ($\mathfrak{gl}_{n}\left(F\right)$). The group/algebra is then pointwise $F$-stable if for any $\varepsilon>0$, there exists a $\delta>0$ such that for any $\delta$-representation $\varphi:S\to GL_{n}\left(F\right)$ ($\mathfrak{gl}_{n}\left(F\right)$), there exists a representation $\psi:S\to GL_{n}\left(F\right)$ ($\mathfrak{gl}_{n}\left(F\right)$) such that $rk\left(\varphi\left(s\right)-\psi\left(s\right)\right)\leq\varepsilon,\;\forall s\in S$. Flexible stability is defined analogously to uniform stability. We refer to [3] for a precise definition and examples, and note that the term stability there corresponds to what we call flexible stability in this paper.
Some form of pointwise $\mathbb{C}$-stability of $\mathbb{Z}^{k}$ were established by Elek and Grabowski in [7], where the $\varepsilon$-representations take values in unitary/self-adjoint matrices. This can be thought of as $k$ matrices that pairwise almost commute, and pointwise stability means that they are close to $k$ matrices that truly commute.
Bauer, Blacher and Greenfeld [3] have studied pointwise $\mathbb{C}$-stability of general polynomial equations. This includes finitely presented groups, associative algebras and Lie algebras. They have proved that finite-dimensional associative algebras are flexibly pointwise stable, but in general are not strictly pointwise stable. Moreover they have proved that finite-dimensional Lie algebras are typically not flexibly stable: If $\mathfrak{g}$ is a finite dimensional complex Lie algebra that is not solvable or has a nonabelian nilpotent quotient, then it is not flexibly $\mathbb{C}$-stable ([3, Theorem 8.5]). Thus any nonabelian finite-dimensional complex Lie algebra is not flexibly $\mathbb{C}$-stable, and the abelian case is open (and is equivalent to pointwise stability of $\mathbb{Z}^{k}$ as in [7]). Although [3] deals with pointwise stability whereas in this paper we study uniform stability, for finite dimensional associative/Lie algebras these notions can be shown to be equivalent, see Section 2.1.
2. Preliminaries and first observations
2.1. Uniform Stability
Let $\Gamma$ be a group and let $\mathcal{G}=\left\{H_{i}\right\}_{i\in I}$ be a family of groups. Let =$\mathcal{F}=\left\{\varphi:\Gamma\to H\;|\;H\in\mathcal{G}\right\}$ be the set of functions from $\Gamma$ to groups in $\mathcal{G}$, a defect function $def:\mathcal{F}\to\mathbb{R}_{\geq 0}$ and a distance function $d:\mathcal{F}\times\mathcal{F}\to\left[0,\infty\right]$. We say that $\Gamma$ is $\left(\mathcal{G},def,d\right)$-stable if for any $\varepsilon>0$ there exists a $\delta>0$ such that for any function $\varphi\in\mathcal{F}$ with $def\left(\varphi\right)\leq\delta$ there exists a homomorphism $\psi\in\mathcal{F}$ such that $d\left(\varphi,\psi\right)\leq\varepsilon$.
In our context, the family $\mathcal{G}$ will be all invertible matrices over $F$, $\mathcal{GL}\left(F\right)=\left\{GL_{n}\left(F\right)\right\}_{n\in\mathbb{N}}$, and the the (uniform) defect function will be
$def\left(\varphi\right)=sup\,\left\{rk\left(\varphi\left(\gamma_{1}\right)\varphi\left(\gamma_{2}\right)-\varphi\left(\gamma_{1}\gamma_{2}\right)\right)\;|\;\gamma_{1},\gamma_{2}\in\Gamma\right\}.$
The distance $d$ with respect to strict uniform stability will be
$d^{strict}\left(\varphi,\psi\right)=\begin{cases}\underset{\gamma\in\Gamma}{sup}\;rk\left(\varphi\left(\gamma\right)-\psi\left(\gamma\right)\right)&\varphi,\psi:\Gamma\to GL_{m}\left(F\right)\\ \infty&else\end{cases},$
so that the codomains of $\varphi$ and $\psi$ must coincide. In the context of flexible stability, we allow taking the distance between maps with different codomains. We use the notation used in [7] and [3]: For $A\in M_{n}\left(F\right)$, let $\widehat{A}$ be the operator on $\mathbb{C}^{\oplus\mathbb{N}}$ which acts as $A$ on the first $n$ basis vectors and 00 otherwise. Thus we may think of $\widehat{A}$ as an infinite matrix that vanishes everywhere except for the submatrix $\widehat{A}_{\left[n\right],\left[n\right]}$ where it equals $A$. For $A\in GL_{n}\left(F\right)$ and $B\in GL_{m}\left(F\right)$, define
$rk\left(A-B\right)\vcentcolon=\frac{dim\left(im\left(\widehat{A}-\widehat{B}\right)\right)}{min\left(n,m\right)},$
And define the flexible distance of $\varphi,\psi\in\mathcal{F}$ by
$d^{flex}\left(\varphi,\psi\right)=\underset{\gamma\in\Gamma}{sup}\,rk\left(\varphi\left(\gamma\right)-\psi\left(\gamma\right)\right).$
We note that since we require the target matrices to be nonsingular, we have
$dim\left(im\left(\widehat{A}-\widehat{B}\right)\right)\geq\left|dim\left(im\left(\widehat{A}\right)\right)-dim\left(im\left(\widehat{B}\right)\right)\right|=\left|n-m\right|,$
so that $rk\left(A-B\right)$ can be small only if $\frac{\left|n-m\right|}{min\left(n,m\right)}$ is small. Thus, in flexible stability we allow an increase of the target dimension, but this increase must be small relative to the dimension.
The group $\Gamma$ is strictly uniformly $F$-stable if it is $\left(\mathcal{GL}\left(F\right),def,d^{strict}\right)$-stable and flexibly uniformly $F$-stable if it is $\left(\mathcal{GL}\left(F\right),def,d^{flex}\right)$-stable.
[…]
Uniform stability of a Lie algebra $\mathfrak{g}$ over a field $K$ can be defined similarly to stability of groups, where $\mathcal{G}=\left\{\mathfrak{h}_{i}\right\}_{i\in I}$ is now a family of Lie algebras over a field $F$ extending $K$, and $\mathcal{F}$ are all $F$-linear maps $\phi:\mathfrak{g}\to\mathfrak{h}_{i}$. The defect of $\phi\in\mathcal{F}$ is then
$def\left(\phi\right)=sup\,\left\{rk\left(\left[\phi\left(x\right),\phi\left(y\right)\right]-\phi\left(\left[x,y\right]\right)\right)\;|\;x,y\in\mathfrak{g}\right\}.$
Strict/flexible uniform $F$- stability of $\mathfrak{g}$ is then taken with respect to $\mathcal{G}=\left\{\mathfrak{gl}_{n}\left(F\right)\right\}_{n\in\mathbb{N}}$ and $d^{strict},d^{flex}$ are defined similarly to the case of groups.
[…]
It is worth noting that one can similarly show equivalence between pointwise and uniform stability of finite dimensional associative algebras, although we do not discuss associative algebras in this paper.
Like in the case of groups, flexible stability of finite dimensional Lie algebras can also be reduced to a small increase of the dimension; By Lemma 3.2 in [3], if $\mathfrak{g}$ is an $m$-dimensional Lie algebra, $\varphi:\mathfrak{g}\to\mathfrak{gl}_{n}\left(F\right)$ is a $\delta$-representation and $\psi:\mathfrak{g}\to\mathfrak{gl}_{N}\left(F\right)$ a representation such that $n\leq N$ and $d^{flex}\left(\varphi,\psi\right)\leq\varepsilon$ then we may assume that $N\leq\left(1+\varepsilon m\right)n$.
2.2. Compressions and strict vs. flexible stability
Fix a subspace $U\leq F^{n}$ of dimension $k$, a basis $B=\left\{u_{1},\ldots,u_{k}\right\}$ for $U$, $\iota:U\to F^{n}$ the inclusion and $p:F^{n}\twoheadrightarrow U$ a projection, i.e. a linear map satisfying $p^{2}=p$ and $im\,p=U$. For any $M\in M_{n}\left(F\right)$, we denote by
$\left[M\right]_{U,B,p}\vcentcolon=p\circ M\circ\iota\in M_{k}\left(F\right)$
the matrix representing the linear map $p\circ M\circ\iota\in End\left(U\right)$ in the basis $B$, and call it the compression of $M$ associated to $\left(U,B,p\right)$. We will often omit $B$ and $p$ from the notation writing $\left[M\right]_{U}$, where a choice of a basis and a projection are implicit.
Lemma 2.1.
  1. (1)
    Let $U\leq F^{n}$ be a subspace of dimension $k$, and $\left[M\right]_{U,B,p}$ the compression associated to $U$ with a basis $B$ and a projection $p$. Then we have
    $rank\left(\left[M\right]_{U,B,p}\right)\geq rank\left(M\right)-2\left(n-k\right)\quad\forall M\in M_{n}\left(F\right),$
    and for any $M_{1},M_{2}\in M_{n}\left(F\right)$
    $rank\left(\left[M_{1}M_{2}\right]_{U,B,p}-\left[M_{1}\right]_{U,B,p}\left[M_{2}\right]_{U,B,p}\right)\leq n-k.$
  2. (2)
    Let $U_{1},U_{2}\leq F^{n}$ be subspaces of dimension $k$, $\iota_{i}:U_{i}\hookrightarrow F^{n}$, $p_{i}:F^{n}\twoheadrightarrow U_{i}$ inclusions and projections, and $B_{i}$ bases for $U_{i}$, respectively. Then for some $A\in GL_{k}\left(F\right)$,
    $rank\left(\left[M\right]_{U_{1},B_{1},p_{1}}-A\left[M\right]_{U_{2},B_{2},p_{2}}A^{-1}\right)\leq 4\left(n-k\right),\quad\forall M\in M_{n}\left(F\right).$
Proof.
We have by the triangle inequality
$rank\left(\left[M\right]_{U}\right)$ $=rank\left(p\circ M\circ\iota\right)=rank\left(\iota\circ p\circ M\circ\iota\circ p\right)$
$=rank\left(M-M\left(I_{n}-\iota\circ p\right)-\left(I_{n}-\iota\circ p\right)M\circ\iota\circ p\right)$
$\geq rank\left(M\right)-2rank\left(I_{n}-\iota\circ p\right)=rank\left(M\right)-2\left(n-k\right).$
For $M_{1},M_{2}\in M_{n}\left(F\right)$,
$rank\left(\left[M_{1}M_{2}\right]_{U}-\left[M_{2}\right]_{U}\left[M_{2}\right]_{U}\right)$ $=rank\left(p\circ\left(M_{1}M_{2}-M_{1}\left(\iota\circ p\right)M_{2}\right)\circ\iota\right)$
$\leq rank\left(M_{1}M_{2}-M_{1}\left(\iota\circ p\right)M_{2}\right)$
$=rank\left(M_{1}\left(I_{n}-\iota\circ p\right)M_{2}\right)$
$\leq rank\left(I_{n}-\iota\circ p\right)=n-k,$
proving 1. For 2, let $W=U_{1}\cap U_{2}$, and let $B_{W}=\left\{w_{1},\ldots,w_{m}\right\}$ be a basis for $W$. By conjugating the whole expression
$\left[M\right]_{U_{1},B_{1},p_{1}}-A\left[M\right]_{U_{2},B_{2},p_{2}}A^{-1}$
by some matrix in $GL_{k}\left(F\right)$, we may assume that the basis $B_{1}$ is some completion of $B_{W}$ to a basis of $U_{1}$. Now by choosing $A\in GL_{k}\left(F\right)$ we may further assume that the basis $B_{2}$ is given by some completion of $B_{W}$ to a basis of $U_{2}$ and that the expression is given by
$\left[M\right]_{U_{1},B_{1},p_{1}}-\left[M\right]_{U_{2},B_{2},p_{2}}.$
Now for $w_{i}\in W$,
$\left(p_{1}\circ M\circ\iota_{1}-p_{2}\circ M\circ\iota_{2}\right)w_{i}=\left(p_{1}-p_{2}\right)Mw_{i},$
thus the restriction of $\left[M\right]_{U_{1},B_{1},p_{1}}-\left[M\right]_{U_{2},B_{2},p_{2}}$ to the subspace $W$ that has dimension at least $2k-n$ has rank at most $rank\left(p_{1}-p_{2}\right)\leq n-dimW\leq 2\left(n-k\right)$. It follows that
$rank\left(\left[M\right]_{U_{1},B_{1},p_{1}}-\left[M\right]_{U_{2},B_{2},p_{2}}\right)\leq 2\left(n-k\right)+n-\left(2k-n\right)=4\left(n-k\right).$
Lemma 2.1 implies that a compression of an almost representation (and in particular, of a representation) onto a large subspace is again an almost representation (and that compressions of nonsingular matrices have almost full rank), and that up to low rank update and conjugation, all $k$-dimensional compressions of almost representations are equivalent. In particular, if all representations have enough subrepresentations, this implies that flexible stability is equivalent to strict stability.
Corollary 2.2.
Let $F$ be an algebraically closed field of characteristic zero. Then a solvable Lie algebra $\mathfrak{g}$ over $K\leq F$ is flexibly uniformly $F$-stable if and only if it is strictly uniformly $F$-stable.
Proof.
Strict stability clearly implies flexible stability. Assume $\mathfrak{g}$ is flexibly stable. Then for any $\delta>0$ there is an $\varepsilon>0$ such that for any $\delta$-representation $\varphi:\mathfrak{g}\to\mathfrak{gl}_{n}\left(F\right)$ there is a representation $\psi:\mathfrak{g}\to\mathfrak{gl}_{N}\left(F\right)$ such that
$\underset{x\in\mathfrak{g}}{sup}\,rk\left(\widehat{\varphi\left(x\right)}-\widehat{\psi\left(x\right)}\right)\leq\varepsilon,$
and we may assume $N\leq\left(1+\varepsilon m\right)n$. By Lie’s theorem, there is a complete flag $F^{N}=V_{N}\supset V_{N-1}\supset\cdots\supset V_{0}=0$ of invariant subspaces of $\psi$ with $dim\,V_{i}=i$. Then for any $k$ (and choice of projection and basis $p_{k},B_{k}$), the compression
$\left[\psi\right]_{V_{k}}=p_{V_{k}}\circ\psi\circ\iota_{V_{k}}:\mathfrak{g}\to\mathfrak{gl}_{n}\left(F\right)$
is a representation, and
$\underset{x\in\mathfrak{g}}{sup}\;rk\left(p_{V_{k}}\circ\widehat{\varphi\left(x\right)}\circ\iota_{V_{k}}-\widehat{\left[\psi\right]_{V_{k}}\left(x\right)}\right)\leq\underset{x\in\mathfrak{g}}{sup}\;rk\left(\widehat{\varphi\left(x\right)}-\widehat{\psi\left(x\right)}\right)\leq\varepsilon.$
Thinking of the matrices as living in $\mathfrak{gl}_{N}\left(F\right)$, by definition of $\widehat{\varphi\left(x\right)}$ it has an $n$-dimensional compression that equals $\varphi$, so by Lemma 2 there is a $A\in GL_{n}\left(F\right)$ such that
$rank\left(\varphi\left(x\right)-A\left(p_{V_{n}}\circ\widehat{\varphi\left(x\right)}\circ\iota_{V_{n}}\right)A^{-1}\right)\leq 4\left(N-n\right)\leq 4m\varepsilon n\quad\forall x\in\mathfrak{g}$
$\implies\underset{x\in\mathfrak{g}}{sup}\;rk\left(\varphi\left(x\right)-A\left[\psi\right]_{V_{n}}A^{-1}\left(x\right)\right)$
$\leq\underset{x\in\mathfrak{g}}{sup}\;rk\left(\varphi\left(x\right)-A\left(p_{V_{n}}\circ\widehat{\varphi\left(x\right)}\circ\iota_{V_{n}}\right)A^{-1}\right)+rk\left(A\left(p_{V_{n}}\circ\widehat{\varphi\left(x\right)}\circ\iota_{V_{n}}-\widehat{\left[\psi\right]_{V_{n}}\left(x\right)}\right)A^{-1}\right)$
$\leq\left(4m+1\right)\varepsilon,$
where $A\left[\psi\right]_{V_{n}}A^{-1}$ is an $n$-dimensional representation. ∎
In the case of finite dimensional $\mathfrak{g}$, Corollary 2.2 can be helpful only in the abelian case, as we know from [3] that nonabelian solvable Lie algebras are not flexibly stable. We note that stability of abelian Lie algebras are equivalent to pointwise stability of free abelian groups, since from bi-invariance
$rk\left(ABA^{-1}B^{-1}-I_{n}\right)=rk\left(AB-BA\right),$
so the same corollary holds for free abelian groups. The proof of Corollary 2.2 is easily seen to hold also for uniform stability of free abelian groups. However, currently it is not even known whether $\mathbb{Z}$ is uniformly $F$-stable for any field $F$. By contrast, finite groups were proven to be $F$-stable in [12, Section 6], for any field $F$. Pointwise and uniform stability for finite groups are easily seen to be equivalent, similar to the case of finite dimensional Lie algebras.
Remark 2.3.
We have defined uniform stability only for individual groups/algebras. However, it is also common to consider uniform stability for families of groups, such as amenable groups in Kazhdan’s theorem, where the dependence of $\delta$ on $\varepsilon$ is uniform for the entire family. The uniform rank metric stability of finite groups in [12, Section 6] depends on the size of the group, and so it does not prove uniform stability for the class of finite groups.
2.3. Verma modules and central characters
Let $\mathfrak{g}$ be a semisimple Lie algebra over $\mathbb{C}$, with a Cartan subalgebra $\mathfrak{h}$ of dimension $\ell$, the rank of $\mathfrak{g}$. Denote by $\Phi\subset\mathfrak{h}^{*}$ the root system of $\mathfrak{g}$ relative to $\mathfrak{h}$. To each root $\alpha$ denote by $\mathfrak{g}_{\alpha}$ the corresponding root space, and fix a set of positive roots $\Phi^{+}\subset\Phi$ and a set of simple roots $\Delta\subset\Phi^{+}$ having $\ell$ elements. This defines a Cartan decomposition
$\mathfrak{g}=\mathfrak{n}^{-}\oplus\mathfrak{h}\oplus\mathfrak{n}^{+},$
where
$\mathfrak{n}^{+}\vcentcolon=\bigoplus_{\alpha>0}\mathfrak{g}_{\alpha}\;,\;\mathfrak{n}^{-}\vcentcolon=\bigoplus_{\alpha<0}\mathfrak{g}_{\alpha},$
with corresponding Borel subalgebras
$\mathfrak{b}^{+}=\mathfrak{h}\oplus\mathfrak{n}^{+}\;,\;\mathfrak{b}^{-}=\mathfrak{h}\oplus\mathfrak{n}^{-}$
Let $\Lambda_{r}$ be the root lattice and $\Lambda$ be the integral weight lattice in $\Phi$, and $\Lambda^{+}\subseteq\Lambda$ the dominant integral weights. Let $\rho$ be the half sum of the positive roots, also called the Weyl vector. Denote by $W$ the (finite) Weyl group of $\Phi$, generated by the reflections $s_{\alpha}$.
There is a standard basis for $\mathfrak{g}$ consisting of $y_{1},\ldots,y_{m}\in\mathfrak{n}^{-}$, $h_{1},\ldots,h_{\ell}\in\mathfrak{h}$ and $x_{1}\ldots,x_{m}\in\mathfrak{n}^{+}$, such that $x_{i}\in\mathfrak{g}_{\alpha_{i}}$ for positive roots $\alpha_{i}\in\Phi^{+}$, $y_{i}\in\mathfrak{g}_{-\alpha_{i}}$, and $h_{1}=h_{\alpha_{i}}$ for simple roots $\alpha_{i}\in\Delta$. According to the Poincaré-Birkhoff-Witt (PBW) theorem, there is a corresponding basis for the universal enveloping algebra $U\left(\mathfrak{g}\right)$ consisting of monomials
$y_{1}^{r_{1}}\cdots y_{m}^{r_{m}}h_{1}^{s_{1}}\cdots h_{\ell}^{s_{\ell}}x_{1}^{t_{1}}\cdots x_{m}^{t_{m}}$
A $U\left(\mathfrak{g}\right)$-module is a highest weight module of weight $\lambda\in\mathfrak{h}^{*}$ if there is a vector $v_{0}\in M$ of weight $\lambda$ such that $\mathfrak{n}^{+}\cdot v_{0}=0$ and $M=U\left(\mathfrak{g}\right)\cdot v_{0}$. By the PBW theorem, such a module satisfies $M=U\left(\mathfrak{n}^{-}\right)\cdot v_{0}$. A universal highest weight module of weight $\lambda$ is called a Verma module and is denoted by $M\left(\lambda\right)$. One may construct it by induction from $\mathfrak{b}$: Since $\mathfrak{b}/\mathfrak{n}$ is isomorphic to $\mathfrak{h}$, there is a $1$-dimensional representation $\mathbb{C}_{\lambda}$ of $\mathfrak{b}$ with trivial $\mathfrak{n}$-action and $\mathfrak{h}$ acting as $\lambda$. By PBW, $U\left(\mathfrak{b}\right)\leq U\left(\mathfrak{g}\right)$ is a subalgebra and thus acts on $U\left(\mathfrak{g}\right)$ by left multiplication.
Set
$M\left(\lambda\right)\vcentcolon=Ind_{\mathfrak{b}}^{\mathfrak{g}}\mathbb{C}_{\lambda}=U\left(\mathfrak{g}\right)\otimes_{U\left(\mathfrak{b}\right)}\mathbb{C}_{\lambda}$
Which has a structure of a left $U\left(\mathfrak{g}\right)$-module with highest weight vector $1\otimes 1$. Again by the PBW theorem, $U\left(\mathfrak{g}\right)\cong U\left(\mathfrak{n}^{-}\right)\otimes U\left(\mathfrak{b}\right)$ and thus $M\left(\lambda\right)\cong U\left(\mathfrak{n}^{-}\right)\otimes\mathbb{C}_{\lambda}$ as a left $U\left(\mathfrak{n}^{-}\right)$-module, and is therefore a free $U\left(\mathfrak{n}^{-}\right)$-module of rank $1$. Every highest weight module for $\lambda$ is a quotient of $M\left(\lambda\right)$, and each $M\left(\lambda\right)$ has a unique simple quotient $L\left(\lambda\right)$ which is finite-dimensional if and only if $\lambda\in\Lambda^{+}$.
We define the shifted action of $W$ on $\mathfrak{h}^{*}$ by $w\cdot\lambda=w\left(\lambda+\rho\right)-\rho$, and we say that $\lambda,\mu\in\mathfrak{h}^{*}$ are $W$-linked if $\mu=w\cdot\lambda$ for some $w\in W$.
The center $Z\left(U\left(\mathfrak{g}\right)\right)$ acts on a highest weight module $M$ of weight $\lambda$ via scalars:
$z\cdot v=\chi_{\lambda}\left(z\right)v\quad\forall v\in M$
Thus $\chi_{\lambda}:Z\left(U\left(\mathfrak{g}\right)\right)\to\mathbb{C}$ is an algebra homomorphism that we call the central character associated with $\lambda$. A fundemental theorem on central characters is due to Harish-Chandra [10, Theorem 1.10]:
Theorem.
(Harish-Chandra isomorphism)
  1. (1)
    There is an isomorphism of algebras $Z\left(U\left(\mathfrak{g}\right)\right)\cong S\left(\mathfrak{h}\right)^{W}$ where $S\left(\mathfrak{h}\right)^{W}$ is the algebra of $W$-invariant elements of the symmetric algebra of $\mathfrak{h}$.
  2. (2)
    For $\lambda,\mu\in\mathfrak{h}^{*}$, $\chi_{\lambda}=\chi_{\mu}$ if and only if $\mu$ and $\lambda$ are $W$-linked.
  3. (3)
    Every central character $\chi:Z\left(U\left(\mathfrak{g}\right)\right)\to\mathbb{C}$ is of the form $\chi_{\lambda}$ for some $\lambda\in\mathfrak{h}^{*}$.
The algebra $S\left(\mathfrak{h}\right)^{W}$ of invariants is isomorphic to a polynomial algebra in $\ell$ variables, and therefore $Z\left(U\left(\mathfrak{g}\right)\right)$ is (freely) generated as an algebra by $\ell$ elements.
2.4. Lattices in semisimple Lie groups
Let $G$ be a semisimple Lie group. A discrete subgroup $\Gamma\leq G$ is a lattice if $G/\Gamma$ has finite volume (i.e. there is a $\sigma$-finite, $G$-invariant Borel measure $\nu$ on $G/\Gamma$). Lattices possess various rigidity properties, one of the most fundemental properties is the Borel Density Theorem. The following formulation is taken from [14, Corollary 4.5.6].
Theorem (Borel Density).
Let $G\leq SL_{r}\left(\mathbb{R}\right)$ be a connected linear semisimple Lie group and $\Gamma\leq G$ a lattice. If $\Gamma$ projects densely onto the maximal compact factor of $G$, then $\Gamma$ is Zariski dense in $G$. That is, if $Q\in\mathbb{R}\left[x_{1,1},\ldots,x_{r,r}\right]$ is a polynomial function on $M_{r}\left(\mathbb{R}\right)$ such that $Q\left(\Gamma\right)=0$, then $Q\left(G\right)=0$.
A remarkable theorem on lattices in higher rank is Margulis Superrigidity, which states that in higher rank $G$ and $\Gamma\leq G$ an irreducible lattice, essentially all finite dimensional representations of $\Gamma$ can be extended to rational representations of $G$ up to some bounded error. For noncocompact $\Gamma$, any representation can be extended to a rational representation of $G$ on some finite index subgroup. For cocompact lattices this is not necessarily true - there may be unitary representations that do not factor through a finite quotient. But we know from Margulis arithmeticity that $\Gamma$ is an arithmetic subgroup of $G$, and therefore we may add compact factors to $G$ so that $\Gamma$ is commensurable to the integral points of $G^{\prime}=G\times K$ (these compact factors are in fact Galois twistings corresponding to the number field defining the arithmetic group $\Gamma$, see [14, Section 5.5]). Then it is true that finite dimensional representations of $\Gamma$ extend on some finite index subgroup to a representation of $G^{\prime}$ (see [14, Section 16.4]).
Theorem (Margulis Superrigidity).
Let $G$ be a connected, algebraically simply connected semisimple Lie group of rank $\geq 2$ without compact factors that is defined over $\mathbb{Q}$ and let $\Gamma\leq G$ be an irreducible lattice. Then there is a compact semisimple group $K$, $G^{\prime}=G\times K$ where $\Gamma\leq G^{\prime}$ is an irreducible lattice projecting densely onto $K$, such that for any representation $\varphi:\Gamma\to GL_{n}\left(\mathbb{C}\right)$ there is a finite index subgroup $\tilde{\Gamma}\leq\Gamma$ and a rational representation $\Phi:G^{\prime}\to GL_{n}\left(\mathbb{C}\right)$ such that $\Phi\mid_{\tilde{\Gamma}}=\varphi$.
3. Semisimple Lie algebras and Lie groups
3.1. Non-stability of semisimple Lie algebras
Let $\mathfrak{g}$ be a finite dimensional complex semisimple Lie algebra. The following theorem shows that although $\mathfrak{g}$ has only finitely many $n$-dimensional representations, it has uncountably many $\varepsilon_{n}$-representations (for some $\varepsilon_{n}\underset{n\to\infty}{\longrightarrow}0$) that are nonequivalent:
Theorem 3.1.
Let $\mathfrak{g}$ be a complex semisimple Lie algebra of rank $\ell$ and $m=\left|\Phi^{+}\right|$ the number of positive roots. Then there is a family
$\left\{\varphi_{n}^{\lambda}:\mathfrak{g}\to\mathfrak{gl}_{k_{n}}\left(\mathbb{C}\right)\;|\;\lambda\in\mathfrak{h}^{*}\right\}$
of $\varepsilon_{n}$-representations, where $\varepsilon_{n}=\frac{2m^{2}}{n}$ and $k_{n}=\binom{n+m}{m}-1$, satisfying the following properties:
  1. (1)
    The vector space $\mathbb{C}^{k_{n}}=V=\bigoplus_{\mu\in\mathfrak{h}^{*}}V_{\mu}$ is a sum of weight spaces, and there is a vector $v_{0}$ of weight $\lambda$ such that $\varphi_{n}^{\lambda}\left(\mathfrak{n}^{+}\right)\cdot v_{0}=0$ and $\varphi_{n}^{\lambda}\left(\mathfrak{n}^{-}\right)=V$.
  2. (2)
[…]
    $rk\left(\tilde{\varphi}_{n}^{\lambda}\left(z_{i}\right)-\chi_{\lambda}\left(z_{i}\right)I_{k_{n}}\right)\leq\left(R+1\right)\varepsilon_{n}$
  3. (3)
    For any $\lambda,\mu\in\mathfrak{h}^{*}$that are not $W$-linked,
    $d^{strict}\left(\varphi_{n}^{\lambda},A\varphi_{n}^{\mu}A^{-1}\right)$ $\geq\left(\binom{R+\ell}{\ell}R\right)^{-1}\left(1-2\left(R+1\right)\varepsilon_{n}\right)$
    $\underset{n\to\infty}{\longrightarrow}\left(\binom{R+\ell}{\ell}R\right)^{-1}$
    For any $A\in GL_{k_{n}}\left(\mathbb{C}\right)$.
  4. (4)
    For any $\lambda\in\mathfrak{h}^{*}$ and any representation $\psi$ of $\mathfrak{g}$ such that $\lambda$ is not $W$-linked to any $\mu$ appearing as a highest weight of a subrepresentation of $\psi$,
    $d^{flex}\left(\varphi_{n}^{\lambda},\psi\right)$ $\geq\left(\binom{R+\ell}{\ell}R\ell\right)^{-1}\left(1-\ell\left(R+m+1\right)\varepsilon_{n}\right)$
    $\underset{n\to\infty}{\longrightarrow}\left(\binom{R+\ell}{\ell}R\ell\right)^{-1}$
    In particular, if $\lambda\in\mathfrak{h}^{*}$ is not $W$-linked to any $\mu\in\Lambda^{+}$, then $\varphi_{n}^{\lambda}$ is bounded from representations of $\mathfrak{g}$, and thus $\mathfrak{g}$ is not flexibly $\mathbb{C}$-stable.
The maps $\varphi_{n}^{\lambda}$ are constructed via finite dimensional compressions of infinite dimensional weight representations of $\mathfrak{g}$ onto subspaces that are “almost invariant”. Looking at weight diagrams of such infinite dimensional representations, it is apparent that $\mathfrak{g}$ acts “locally” on the weight spaces, so that if one takes a large ball in $\mathfrak{h^{*}}$, the relations of $\mathfrak{g}$ are satisfied on the interior and not on the boundary, which can be shown to be small with respect to the whole subspace. One can instead note that $U\left(\mathfrak{g}\right)$ has polynomial growth and use this to construct the almost invariant subspaces taking polynomials up to a given degree in quotients of the $\mathfrak{g}$-module $U\left(\mathfrak{g}\right)$.
Note that the non-stability can already be deduced from (3) using cardinality considerations, however, (4) shows that the non-stability can be witnessed by any weight that is not $W$-linked to dominant integral weights.
Proof.
We construct $\varphi_{n}^{\lambda}$ by compressing the Verma module $M\left(\lambda\right)$ to a finite dimensional subspace. Recall that $M\left(\lambda\right)$ is free as a $U\left(\mathfrak{n}^{-}\right)$-module, and therefore has a basis of the form
$M\left(\lambda\right)=span\left\{y_{1}^{r_{1}}\cdots y_{m}^{r_{m}}\cdot v_{0}\;|\;r_{i}\in\mathbb{Z}_{\geq 0}\right\},$
where $y_{1},\ldots,y_{m}$ are the negative root spaces and $v_{0}$ the highest weight vector. For $n\in\mathbb{N}$, Denote by
$D_{n}=span\left\{y_{1}^{r_{1}}\cdots y_{m}^{r_{m}}\cdot v_{0}\;|\;\sum_{i=1}^{m}r_{i}\leq n\right\}$
the subspace spanned by all monomials in $U\left(\mathfrak{n}^{-}\right)$ of degree $\leq n$ acting on $v_{0}$. A standard partition argument shows that there are $\binom{n+m-1}{m-1}$ such combinations for fixed $a$, and
$dim\left(D_{n}\right)=\sum_{j=1}^{n}\binom{j+m-1}{m-1}=\binom{n+m}{m}-1.$
Define the action of $\mathfrak{g}$ on $D_{n}$ by
$\left(\varphi_{n}^{\lambda}\left(x\right)\right)\left(y_{1}^{r_{1}}\cdots y_{m}^{r_{m}}\cdot v_{0}\right)=\begin{cases}xy_{1}^{r_{1}}\cdots y_{m}^{r_{m}}\cdot v_{0}&xy_{1}^{r_{1}}\cdots y_{m}^{r_{m}}\cdot v_{0}\in D_{n}\\ 0&else\end{cases}.$
We claim that the image of this action of $\mathfrak{g}$ on $D_{n-1}$ is contained in $D_{n}$; For $\mathfrak{n}^{-}$, after reordering using a PBW basis for $\mathfrak{n}^{-}$ we have a sum of monomials of degree at most $n$. The action of $\mathfrak{h}$ is by scalars and therefore any $D_{k}$ is $\mathfrak{h}$-invariant. We prove by induction on $n$ that $\mathfrak{n}^{+}\cdot D_{n}\subseteq D_{n}$: When $n=0$ this is obvious. Assume it holds for $n-1$, then for $x\in\mathfrak{g}_{\beta}$ ($\beta>0$) we have
$xy_{1}^{r_{1}}y_{2}^{r_{2}}\cdots y_{m}^{r_{m}}\cdot v_{0}=y_{1}xy_{1}^{r_{1}-1}y_{2}^{r_{2}}\cdots y_{m}^{r_{m}}\cdot v_{0}+zy_{1}^{r_{1}-1}y_{2}^{r_{2}}\cdots y_{m}^{r_{m}}\cdot v_{0}$
for some $z\in\mathfrak{g}_{-\alpha_{1}+\beta}$. By the induction hypothesis, $xy_{1}^{r_{1}-1}y_{2}^{r_{2}}\cdots y_{m}^{r_{m}}\cdot v_{0}\in D_{n-1}\implies y_{1}xy_{1}^{r_{1}-1}y_{2}^{r_{2}}\cdots y_{m}^{r_{m}}\cdot v_{0}\in D_{n}$. Moreover, if $\beta-\alpha_{1}>0$ then again use the induction hypothesis for $zy_{1}^{r_{1}-1}y_{2}^{r_{2}}\cdots y_{m}^{r_{m}}\cdot v_{0}$, and if $\beta-\alpha_{1}\leq 0$ then $z\in\mathfrak{n}^{-}$ or $z\in\mathfrak{h}$ and in any case $zy_{1}^{r_{1}-1}y_{2}^{r_{2}}\cdots y_{m}^{r_{m}}\cdot v_{0}\in D_{n}$.
Now since $\varphi_{n}^{\lambda}\left(\mathfrak{g}\right)\cdot D_{n-1}\subseteq D_{n}$, we have $\left[\varphi_{n}^{\lambda}\left(x\right),\varphi_{n}^{\lambda}\left(y\right)\right]\cdot D_{n-2}\subseteq D_{n}$ for any $x,y\in\mathfrak{g}$. Since $\varphi_{n}^{\lambda}$ acts as $M\left(\lambda\right)$ when restriced to $D_{n-2}$, we have $\left(\varphi_{n}^{\lambda}\left(\left[x,y\right]\right)-\left[\varphi_{n}^{\lambda}\left(x\right),\varphi_{n}^{\lambda}\left(y\right)\right]\right)\cdot D_{n-2}=0$. Thus
$rank\left(\varphi_{n}^{\lambda}\left(\left[x,y\right]\right)-\left[\varphi_{n}^{\lambda}\left(x\right),\varphi_{n}^{\lambda}\left(y\right)\right]\right)$ $\leq dim\,D_{n}-dim\,D_{n-2}=\binom{n+m-1}{n}+\binom{n+m-2}{n}$
$\implies\varepsilon_{n}\leq\frac{\binom{n+m-1}{n}+\binom{n+m-2}{m-1}}{\binom{n+m}{m}}$ $=\frac{m}{n+m}\left(1+\frac{mn}{\left(n+m-1\right)}\right)\leq\frac{m}{n+m}\left(1+m\right)<\frac{2m^{2}}{n}.$
All of the generators $z_{1},\ldots,z_{\ell}$ of $Z\left(U\left(\mathfrak{g}\right)\right)$ have degree at most $R$, so when restricted to $D_{n-R}$ they act as the scalars $\chi_{\lambda}\left(z_{i}\right)$. Thus
$rank\left(\tilde{\varphi}_{n}^{\lambda}\left(z_{i}\right)-\chi_{\lambda}\left(z_{i}\right)I_{k_{n}}\right)$ $\leq dimD_{n}-dimD_{n-R}$
$=\binom{n+m}{m}-\binom{n-R+m}{m}$
$\implies rk\left(\tilde{\varphi}_{n}^{\lambda}\left(z_{i}\right)-\chi_{\lambda}\left(z_{i}\right)I_{k_{n}}\right)\leq 1-\frac{\binom{n-R+m}{m}}{\binom{n+m}{m}}\leq 1-\frac{\left(n-R\right)^{m}}{\left(n+m\right)^{m}}.$
Using the inequality $log\left(1+x\right)\geq\frac{x}{1+x}$,
$log\left(\frac{\left(n-R\right)^{m}}{\left(n+m\right)^{m}}\right)=mlog\left(1-\frac{m+R}{n+m}\right)\geq-m\frac{m+R}{n-R},$
so that
$rk\left(\tilde{\varphi}_{n}^{\lambda}\left(z_{i}\right)-\chi_{\lambda}\left(z_{i}\right)I_{k_{n}}\right)$ $\leq 1-exp\left(-m\frac{m+R}{n-R}\right)\leq m\frac{m+R}{n-R}$
$<m\frac{m+2R}{n}<\left(R+1\right)\varepsilon_{n}.$
This proves 2.
For 3, by Harish-Chandra’s theorem $\chi_{\lambda}\neq\chi_{\mu}$ so that they differ on at least one of the generators, say $z_{j}$. Therefore, for any $A\in GL_{k_{n}}\left(\mathbb{C}\right)$,
$rk\left(\tilde{\varphi}_{n}^{\lambda}\left(z_{j}\right)-A\tilde{\varphi}_{n}^{\mu}\left(z_{j}\right)A^{-1}\right)$ $\geq rk\left(\chi_{\lambda}\left(z_{j}\right)-A\chi_{\mu}\left(z_{j}\right)A^{-1}\right)-2\left(R+1\right)\varepsilon_{n}$
$=1-2\left(R+1\right)\varepsilon_{n}.$
Let $\eta=\varphi_{n}^{\lambda}-A\varphi_{n}^{\mu}A^{-1}$ and $\delta=\underset{x\in\mathfrak{g}}{max}\;rk\left(\eta\left(x\right)\right)$. Every $z_{i}$ is a polynomial of degree $\leq R$ in the PBW basis, so it is a sum of at most $\binom{R+\ell}{\ell}-1$ monomials of degree at most $R$. By [3, Lemma 3.3] we therefore have
$rk\left(\tilde{\varphi}_{n}^{\lambda}\left(z_{j}\right)-A\tilde{\varphi}_{n}^{\mu}\left(z_{j}\right)A^{-1}\right)\leq\binom{R+\ell}{\ell}R\delta.$
Combining the above we get
$1-2\left(R+1\right)\varepsilon_{n}\leq\binom{R+\ell}{\ell}R\delta$
$\implies\delta\geq\left(\binom{R+\ell}{\ell}R\right)^{-1}\left(1-2\left(R+1\right)\varepsilon_{n}\right),$
proving 3.
Let $\psi:\mathfrak{g}\to\mathfrak{gl}_{N}\left(\mathbb{C}\right)$ be a representation, where $N\leq\left(1+\varepsilon_{n}m\right)k_{n}$, and suppose that $\lambda\in\mathfrak{h}^{*}$is not $W$-linked to any $\mu\in\mathfrak{h}^{*}$ appearing as a highest weight of a subrepresentation of $\psi$. By complete reducibility, it is a (conjugation of) a direct sum of isotypic components
$\psi=A\left(\bigoplus_{i=1}^{q}\psi_{\mu_{i}}\right)A^{-1},$
where $\psi_{\mu_{i}}$ is the direct sum of copies of the irreducible representation with highest weight $\mu_{i}$ (which must in particular be dominant integral weights). The central character of $\psi$ is therefore
$\chi_{\psi}=A\left(\bigoplus_{i=1}^{q}\chi_{\mu_{i}}I_{d_{i}}\right)A^{-1}.$
Let $\eta=\widehat{\varphi_{n}^{\lambda}}-\widehat{\psi}$ and $\delta=\underset{x\in\mathfrak{g}}{max}\;rk\left(\eta\left(x\right)\right)$. Similar to the calculation above, for each $1\leq i\leq\ell$ we have
$rk\left(\tilde{\psi}\left(z_{i}\right)-\chi_{\lambda}\left(z_{i}\right)I_{k_{n}}\right)$ $\leq rk\left(\tilde{\psi}\left(z_{i}\right)-A\tilde{\varphi}_{n}^{\lambda}\left(z_{j}\right)A^{-1}\right)+rk\left(A\tilde{\varphi}_{n}^{\lambda}\left(z_{j}\right)A^{-1}-\chi_{\lambda}\left(z_{i}\right)I_{k_{n}}\right)$
$\leq\binom{R+\ell}{\ell}R\delta+\left(R+1\right)\varepsilon_{n}.$
Since $N-k_{n}\leq\varepsilon_{n}m$, we have
$rk\left(\tilde{\psi}\left(z_{i}\right)-\chi_{\lambda}\left(z_{i}\right)I_{N}\right)$ $\leq\binom{R+\ell}{\ell}R\delta+\left(R+1\right)\varepsilon_{n}+\left(N-k_{n}\right)$
$\leq\binom{R+\ell}{\ell}R\delta+\left(R+m+1\right)\varepsilon_{n}.$
Thinking of $\widehat{\varphi_{n}^{\lambda}},\psi$ as embedded in $\mathfrak{gl}_{N}\left(\mathbb{C}\right)$, the subspaces $W_{i}=ker\left(\tilde{\psi}\left(z_{i}\right)-\chi_{\lambda}\left(z_{i}\right)I_{N}\right)\leq\mathbb{C}^{N}$ have dimension at least $k_{n}\left(1-\binom{R+\ell}{\ell}R\delta-\left(R+m+1\right)\varepsilon_{n}\right)$ and for any $v\in W$,
$\tilde{\psi}\left(z_{i}\right)v=\chi_{\lambda}\left(z_{i}\right)v,$
hence $\tilde{\psi}\left(z_{i}\right)$ has an eigenspace of dimension at least $k_{n}\left(1-\binom{R+\ell}{\ell}R\delta-\left(R+m+1\right)\varepsilon_{n}\right)$ corresponding to the eigenvalue $\chi_{\lambda}\left(z_{i}\right)$. Suppose the intersection
$W=\bigcap_{i=1}^{\ell}W_{i}$
contains a nonzero vector $0\neq w\in W$, then $w$ satisfies for any $1\leq i\leq\ell$
$\tilde{\psi}\left(z_{i}\right)w=\chi_{\lambda}\left(z_{i}\right)w,$
and in particular it follows that $w$ must be in some isotypic component of $\psi$. This is a contradiction because $\chi_{\lambda}\neq\chi_{\mu_{i}}$ by assumption and Harish-Chandra’s theorem, so $W=0$. But since all $W_{i}$ are subspaces of $\mathbb{C}^{N}$,
$0=dim\,W$ $\geq N-\sum_{i=1}^{\ell}\left(N-dim\,W_{i}\right)\geq N-\ell k_{n}\left(\binom{R+\ell}{\ell}R\delta+\left(R+m+1\right)\varepsilon_{n}\right)$
$\geq N\left(1-\ell\left(\binom{R+\ell}{\ell}R\delta+\left(R+m+1\right)\varepsilon_{n}\right)\right)$
$\implies 1\leq\ell\left(\binom{R+\ell}{\ell}R\delta\right)+\ell\left(R+m+1\right)\varepsilon_{n}$
$\delta\geq\left(\binom{R+\ell}{\ell}R\ell\right)^{-1}\left(1-\ell\left(R+m+1\right)\varepsilon_{n}\right)$
By complexifying real Lie algebras, we can prove similar results for real Lie algebras.
Proposition 3.2.
Let $\mathfrak{g}$ be a finite dimensional real Lie algebra and $\varphi:\mathfrak{g}\to\mathfrak{gl}_{n}\left(\mathbb{C}\right)$ an $\varepsilon$-representation.
Then there is a $4\varepsilon$-representation $\tilde{\varphi}:\mathfrak{g}_{\mathbb{C}}\to\mathfrak{gl}_{n}\left(\mathbb{C}\right)$ extending $\varphi$. Moreover, if $d^{strict}\left(\varphi,\psi\right)=\delta$ then $d^{strict}\left(\tilde{\varphi},\tilde{\psi}\right)\leq 2\delta$ (and similarly for $d^{flex}$).
Proof.
Define $\tilde{\varphi}$ in the obvious way by $\tilde{\varphi}\left(x+iy\right)=\varphi\left(x\right)+i\varphi\left(y\right)$. Then
$\left[\tilde{\varphi}\left(x+iy\right),\tilde{\varphi}\left(w+iv\right)\right]=\left[\varphi\left(x\right),\varphi\left(w\right)\right]-\left[\varphi\left(y\right),\varphi\left(v\right)\right]+i\left(\left[\varphi\left(x\right),\varphi\left(v\right)\right]+\left[\varphi\left(y\right),\varphi\left(w\right)\right]\right)$
$\tilde{\varphi}\left(\left[x+iy,w+iv\right]\right)=\varphi\left(\left[x,w\right]\right)-\varphi\left(\left[y,v\right]\right)+i\left(\varphi\left(\left[x,v\right]\right)+\varphi\left(\left[y,w\right]\right)\right),$
and it follows that $\tilde{\varphi}$ is a $4\varepsilon$-representation.
If $\psi$ is a representation of $\mathfrak{g}$ such that $d^{strict}\left(\varphi,\psi\right)=\delta$ (respectively $d^{flex}\left(\varphi,\psi\right)=\delta$) then for any $x+iy\in\mathfrak{g}_{\mathbb{C}}$,
$rk\left(\tilde{\varphi}\left(x+iy\right)-\tilde{\psi}\left(x+iy\right)\right)\leq rk\left(\varphi\left(x\right)-\psi\left(x\right)\right)+rk\left(i\left(\varphi\left(y\right)-\psi\left(y\right)\right)\right)\leq 2\delta$
Corollary 3.3.
A real Lie algebra $\mathfrak{g}$ is (flexibly) $\mathbb{C}$-stable if and only if $\mathfrak{g}_{\mathbb{C}}$ is (flexibly) $\mathbb{C}$-stable (as a complex Lie algebra). In particular, by Theorem 3.1, real semisimple Lie algebras are not flexibly $\mathbb{C}$-stable.
Proof.
If $\mathfrak{g}$ is (flexibly) stable, then any $\varepsilon$-representation $\tilde{\varphi}$ of $\mathfrak{g}_{\mathbb{C}}$ restricts to an $\varepsilon$-representation $\varphi$ of $\mathfrak{g}$, which satisfies $d^{strict}\left(\varphi,\psi\right)\leq\delta$ (respectively $d^{flex}\left(\varphi,\psi\right)\leq\delta$) for some representation $\psi$. Complexifying both, by Proposition 3.2 $\tilde{\psi}$ is a representation of $\mathfrak{g}_{\mathbb{C}}$ satisfying $d^{strict}\left(\tilde{\psi},\tilde{\varphi}\right)\leq 2\delta$ (respectively $d^{flex}\left(\varphi,\psi\right)\leq 2\delta$).
Conversely, if $\mathfrak{g}_{\mathbb{C}}$ is (flexibly) stable, then an $\frac{\varepsilon}{4}$-representation $\varphi$ of $\mathfrak{g}$ can be extended by Proposition 3.2 to an $\varepsilon$-representation $\tilde{\varphi}$ of $\mathfrak{g}_{\mathbb{C}}$. Hence there is a representation $\tilde{\psi}$ of $\mathfrak{g}_{\mathbb{C}}$ satisfying $d^{strict}\left(\tilde{\varphi},\tilde{\psi}\right)\leq\delta$ (respectively $d^{flex}\left(\tilde{\varphi},\tilde{\psi}\right)\leq\delta$), and restricting to $\mathfrak{g}$ we have a representation $\psi$ of $\mathfrak{g}$ with $d^{strict}\left(\varphi,\psi\right)\leq\delta$ (respectively $d^{flex}\left(\varphi,\psi\right)\leq\delta$). ∎
3.2. Semisimple Lie groups and lattices
One of the motivations to consider stability of Lie algebras is that almost representations arise naturally as differentials of (smooth) almost representations of Lie groups. Unlike matrix norms, the rank metric is in general compatible with differentiation:
Lemma 3.4.
Let $B_{\delta}\left(M\right)\subseteq M_{n}\left(\mathbb{C}\right)$ be the $\delta$-ball around $M$ in $M_{n}\left(\mathbb{C}\right)$,
$B_{\delta}\left(M\right)=\left\{A\in M_{n}\left(\mathbb{C}\right)\;|\;rk\left(A-M\right)\leq\delta\right\},$
and let $\gamma:\left(-1,1\right)\to B_{\delta}\left(M\right)$ be a smooth curve. Then $\frac{d\gamma}{dt}:\left(-1,1\right)\to B_{2\delta}\left(0\right)$, so that $rk\left(\frac{d\gamma}{dt}\mid_{t_{0}}\right)\leq 2\delta$.
Proof.
Being contained in $B_{\delta}\left(M\right)$ is equivalent to the vanishing of all the $\left(k+1\right)$-minors of $\gamma\left(t\right)-M$ for $k=\left\lfloor\delta n\right\rfloor$, and is therefore a closed condition - thus $rank\left(\bullet\right)$ is a lower semicontinuous function. We therefore have
$rank\left(\underset{h\to 0}{lim}\,\frac{\gamma\left(t_{0}+h\right)-\gamma\left(t_{0}\right)}{h}\right)$ $\leq\underset{h\to 0}{liminf}\,rank\left(\frac{\gamma\left(t_{0}+h\right)-\gamma\left(t_{0}\right)}{h}\right)$
[…]
Proposition 3.5.
  • Let $G$ be a Lie group and $\Phi:G\to GL_{n}\left(\mathbb{C}\right)$ a smooth uniform $\varepsilon$-representation. Then $\varphi\vcentcolon=d\Phi\mid_{id}:\mathfrak{g}\to\mathfrak{gl}_{n}\left(\mathbb{C}\right)$ is an $8\varepsilon$-representation.
  • If $\Psi:G\to GL_{N}\left(\mathbb{C}\right)$ is another smooth map, $\psi=d\Psi$ and $d^{strict}\left(\Phi,\Psi\right)\leq\delta$ (respectively $d^{flex}\left(\Phi,\Psi\right)\leq\delta$), then $d^{strict}\left(\varphi,\psi\right)\leq 2\delta$ ($d^{flex}\left(\varphi,\psi\right)\leq 2\delta$).
Proof.
For any $g,h\in G$ we have by the triangle inequality
$rk\left(\Phi\left(ghg^{-1}\right)-\Phi\left(g\right)\Phi\left(h\right)\Phi\left(g\right)^{-1}\right)\leq 2\varepsilon.$
Written differently setting $\Psi_{g}\left(h\right)=ghg^{-1}$,
$rk\left(\Phi\left(\Psi_{g}\left(h\right)\right)-\Psi_{\Phi\left(g\right)}\left(\Phi\left(h\right)\right)\right)\leq 2\varepsilon.$
Taking the differential at $h=id$ we get by Lemma 3.4 that
(3.1) $rk\left(\varphi\left(Ad_{g}\left(x\right)\right)-Ad_{\Phi\left(g\right)}\left(\varphi\left(x\right)\right)\right)\leq 4\varepsilon,$
for any $x\in\mathfrak{g}$. Taking again the differential at $g=id$,
$rk\left(\varphi\left(ad\left(y\right)\left(x\right)\right)-ad\left(\varphi\left(y\right),\varphi\left(x\right)\right)\right)\leq 8\varepsilon.$
Thus,
$rk\left(\varphi\left(\left[y,x\right]\right)-\left[\varphi\left(y\right),\varphi\left(x\right)\right]\right)\leq 8\varepsilon.$
The second assertion follows by applying again Lemma 3.4 and differentiating $\Phi\left(g\right)-\Psi\left(g\right)$. ∎
A natural way to construct almost representations of a Lie group $G$ is by compressing representations as in Lemma 2.1. It is straightforward to check that taking differentials commutes with compressing, so that the differential of a compression of a representation $\Phi$ of $G$ at $id$ will be a compression of the differential $d\Phi$ on $\mathfrak{g}$.
By 4 in Theorem 3.1, the irreducible representation $\psi$ of $\mathfrak{g}$ with highest weight $\mu$ is bounded from any representations of strictly lower dimension, since clearly such a representation cannot contain $\psi$ as a subrepresentation. It therefore follows from Corollary 3.3 and Proposition 3.5,
Corollary 3.6.
Semisimple Lie groups are not strictly $\mathbb{C}$-stable.
We would like to apply this result to lattices in higher rank. Since $rk\left(\bullet\right)$ is a polynomial function on the entries of a matrix, two maps that are close on a Zariski dense subgroup are close on the entire group:
Lemma 3.7.
Let $G\leq SL_{r}\left(\mathbb{R}\right)$ be a linear Lie group and $H\leq G$ a Zariski dense subgroup. Suppose $\Phi,\Psi:G\to GL_{n}\left(\mathbb{C}\right)$ are rational maps satisfying
$\underset{h\in H}{sup}\,rk\left(\Phi\left(h\right)-\Psi\left(h\right)\right)\leq\varepsilon.$
Then
$\underset{g\in G}{sup}\,rk\left(\Phi\left(g\right)-\Psi\left(g\right)\right)\leq\varepsilon.$
Proof.
The condition $rk\left(\Phi\left(g\right)-\Psi\left(g\right)\right)\leq\varepsilon$ is polynomial in the matrix entries, and is equivalent to the vanishing of all $\left(k+1\right)$-minors for $k=\left\lfloor\varepsilon n\right\rfloor$. Since $\Phi,\Psi$ are rational, it follows from Zariski density of $H$ that $\underset{g\in G}{sup}\,rk\left(\Phi\left(g\right)-\Psi\left(g\right)\right)\leq\varepsilon$. ∎
We may now use Borel Density and apply Lemma 3.7 to lattices in semisimple Lie groups.
Corollary 3.8.
Let $G\leq SL_{r}\left(\mathbb{R}\right)$ be a connected linear semisimple Lie group and $\Gamma\leq G$ a lattice that projects densely into the maximal compact factor of $G$. Suppose $\Phi,\Psi:G\to GL_{n}\left(\mathbb{C}\right)$ are rational maps satisfying
$\underset{\gamma\in\Gamma}{sup}\,rk\left(\Phi\left(\gamma\right)-\Psi\left(\gamma\right)\right)\leq\varepsilon.$
Then
$\underset{g\in G}{sup}\,rk\left(\Phi\left(g\right)-\Psi\left(g\right)\right)\leq\varepsilon.$
Margulis Superrigidity then implies that lattices in higher rank are not strictly stable.
proof of Theorem 1.1.
By adding compact factors to $G$, we may assume WLOG that $\Gamma$ is commensurable to $G_{\mathbb{Z}}$, and that it projects densely onto the maximal compact factor of $G$.
A codimension $1$ compression of a highest weight representation $\Phi:G\to GL_{n}\left(\mathbb{C}\right)$ is a $\frac{2}{n-1}$-representation by Lemma 2.1. Let $\varphi:\Gamma\to GL_{n-1}\left(\mathbb{C}\right)$ be the restriction of $\Phi$ to the lattice $\Gamma$. Suppose $\psi:\Gamma\to GL_{n-1}\left(\mathbb{C}\right)$ is a representation such that $d^{strict}\left(\varphi,\psi\right)=\delta$. Then by Margulis Superrigidity, restricting to a finite index subgroup $\bar{\Gamma}\leq\Gamma$, $\psi$ is the restriction of a representation $\Psi$ of $G$. By Corollary 3.8, since $\Phi,\Psi$ are rational and since $\bar{\Gamma}$ is also a lattice that projects densely into the maximal compact factor, we have
$d^{strict}\left(\Phi,\Psi\right)\leq\delta,$
contradicting Corollary 3.6. ∎
3.3. Remarks on lifting almost representations of Lie algebras
It is natural to ask whether the converse of Proposition 3.5 holds, i.e. whether almost representations of real Lie algebras always arise as differentials of almost representations of the corresponding simply connected Lie group. In particular, whether the $\varepsilon_{n}$-representations $\varphi_{n}^{\lambda}$ in Theorem 3.1, restricted to some real Lie algebra $\mathfrak{g}$, arise as differentials of almost representations of the corresponding Lie group $G$. Indeed, a positive answer for any $\varphi_{n}^{\lambda}$ for $\lambda$ such that $\chi_{\lambda}$ is not the central character of any finite dimensional representation, would imply that Lie groups and lattices in higher rank Lie groups are not flexibly $\mathbb{C}$-stable, using the same reasoning as in the proof of Theorem 1.1. We do not know the answer to this question.
However, we do note that there seem to be some difficulties in lifting such almost representations of $\mathfrak{g}$ to those of $G$; Recall that the Weyl group of $G$ is $W=N_{G}\left(\mathfrak{h}\right)/C_{G}\left(\mathfrak{h}\right)$, so by equation 3.1 in the proof of Proposition 3.5, we have
$rk\left(\varphi\left(wxw^{-1}\right)-\Phi\left(w\right)\varphi\left(x\right)\Phi\left(w\right)^{-1}\right)\leq 4\varepsilon,$
for any $x\in\mathfrak{g}$ and $w\in W$. We can interpret this by saying that $\varphi$ has to be almost invariant under the Weyl group action.
Consider the simplest case of $\mathfrak{g}=\mathfrak{sl}_{2}\left(\mathbb{C}\right)$ (whose representations can be seen as the building blocks of representations of all semisimple Lie algebras). Let
$\mathbf{e}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix},\quad\mathbf{f}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix},\quad\mathbf{h}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$
be the standard basis of $\mathfrak{g}$, then the nontrivial element $w\in W$ acts on $\mathfrak{g}$ by
$\mathbf{h}\mapsto-\mathbf{h},\;\mathbf{e}\mapsto\mathbf{-\mathbf{f}},\;\mathbf{f}\mapsto-\mathbf{e},$
which when applied to $L\left(\lambda\right)$ results in the lowest weight module $\bar{L}\left(-\lambda\right)$ with lowest weight $-\lambda$ (see [13, 3]). When $\lambda$ is dominant integral (in this case $\lambda\in\mathbb{Z}_{\geq 0}$), these modules are indeed isomorphic. But if $\lambda\notin\mathbb{Z}_{\geq 0}$, this is not true, but they do have the same central character. If $\varphi_{n}^{\lambda}$ can be lifted to some $\Phi_{n}^{\lambda}$, this means that finite dimensional compressions of $L\left(\lambda\right)$ are arbitrarily close in rank to compressions of (a conjugation of) $\bar{L}\left(-\lambda\right)$. The technique of central characters only helps in separating weight modules with distinct central characters, but in $\mathfrak{sl}_{2}\left(\mathbb{C}\right)$ for every central character $\chi_{\lambda}$ there is a family of weight modules
$\left\{V\left(\xi,\chi_{\lambda}\right)\;|\;\xi\in\mathbb{C}/2\mathbb{Z}\right\}$
having central character $\chi_{\lambda}$ and spectrum $\xi$ (see [13, 3.3]). We do not know whether modules with the same central character but different spectrum can be close to one another, in the sense that their respective finite dimensional compressions are arbitrarily close in the rank metric.
4. Free groups
In [15], Rolli constructed uniform almost representations of nonabelian free groups in the operator norm that are bounded from true representations (see also [5, Section 3]). This inspired a construction by Becker and Chapman [4, Section 5] of uniform almost homomorphisms to permutation groups $S_{n}$, that are bounded from true representations.
We construct similar almost representations in the rank metric inspired by Rolli’s construction. It is interesting to note that it is in some sense opposite to Rolli’s construction; There, the maps $\varphi$ behave locally like representations that are contained in some $\delta$-ball around the identity, but since $GL_{n}\left(\mathbb{C}\right)$ has no small subgroups, it is bounded from true representations where high powers of the generators must be bounded away from the identity. In the case of the rank metric, a finite set of elements close to the identity will generate an algebra that is contained in a small ball around the identity (so in a sense there are “many” small subgroups), but the almost representations that we construct have elements as far as possible from the identity.
Proposition 4.1.
Let $\mathbb{F}_{2}=\left\langle a,b\right\rangle$ and $n\in\mathbb{N}$. Then there is a $\frac{3}{n}$-representation $\varphi:\mathbb{F}_{2}\to GL_{n}\left(F\right)$ such that
$rk\left(\varphi\left(a\right)-I_{n}\right),rk\left(\varphi\left(b\right)-I_{n}\right)\leq\frac{1}{n},$
and such that there is an element $w\in\mathbb{F}_{2}$ satisfying
$rk\left(\varphi\left(w\right)-I_{n}\right)=1.$
Moreover, for any representation $\psi:\mathbb{F}_{2}\to GL_{N}\left(F\right)$, $d^{flex}\left(\varphi,\psi\right)\geq\frac{1}{6}-\frac{1}{6n}$.
Proof.
Let $\delta=\frac{1}{n}$ and let
$B_{\delta}=\left\{A\in GL_{n}\left(F\right)\;|\;rk\left(A-I\right)\leq\delta\right\}$
be the $\delta$-ball around $I$. Let $\tau:\mathbb{Z}\to B_{\delta}$ be a symmetric map so that $\tau\left(k\right)^{-1}=\tau\left(k^{-1}\right)$, $\tau\left(0\right)=I$, and define $\varphi:\mathbb{F}_{2}\to GL_{n}\left(F\right)$ by acting on a reduced word of the form
$w=a^{n_{1}}b^{m_{1}}\cdots a^{n_{k}}b^{m_{k}}$
by
$\varphi\left(w\right)=\tau\left(m_{1}\right)\cdots\tau\left(m_{k}\right).$
Allowing $n_{1}=0$ and $m_{k}=0$ defines $\varphi$ on all reduced words. Now for words $w_{1}=a^{n_{1}}b^{m_{1}}\cdots a^{n_{k}}b^{m_{k}}$, $w_{2}=a^{\ell_{1}}b^{q_{1}}\cdots a^{\ell_{d}}b^{q_{d}}$, if $m_{k}\neq 0$ and $\ell_{1}\neq 0$, there is no cancellation and $\varphi\left(w_{1}w_{2}\right)=\varphi\left(w_{1}\right)\varphi\left(w_{2}\right)$, and the same occurs if $m_{k}=\ell_{1}=0$. If exactly one of $m_{k},\ell_{1}$ equals 00, assume WLOG that $m_{k}=0$ and $\ell_{1}\neq 0$, and the other case is proven similarly. We have
$w_{1}w_{2}=a^{n_{1}}b^{m_{1}}\cdots a^{n_{k-r}}b^{m_{k-r}}ca^{\ell_{r}}b^{q_{r}}\cdots a^{\ell_{d}}b^{q_{d}}$
for $c$ of the form $c=a^{\alpha}b^{\beta}$, which occurs if $n_{k-i+1}=-\ell_{i}$ and $m_{k-i}=-q_{i}$ for $21\leq i\leq r-$ 2. Symmetry of $\tau$ ensures that the corresponding powers cancel so that
$\varphi\left(w_{1}\right)\varphi\left(w_{2}\right)$ $=\tau\left(m_{1}\right)\cdots\tau\left(m_{k}\right)\tau\left(q_{1}\right)\cdots\tau\left(q_{d}\right)$
$=\tau\left(m_{1}\right)\cdots\tau\left(m_{k-r}\right)\tau\left(m_{k-\left(r-1\right)}\right)\tau\left(q_{r-1}\right)\tau\left(q_{r}\right)\cdots\tau\left(q_{d}\right),$
and
$\varphi\left(w_{1}w_{2}\right)=\tau\left(m_{1}\right)\cdots\tau\left(m_{k-r}\right)\tau\left(\beta\right)\tau\left(q_{r}\right)\cdots\tau\left(q_{d}\right)$
$\implies rk\left(\varphi\left(w_{1}w_{2}\right)-\varphi\left(w_{1}\right)\varphi\left(w_{2}\right)\right)=rk\left(\tau\left(m_{k-\left(r-1\right)}\right)\tau\left(q_{r-1}\right)-\tau\left(\beta\right)\right)\leq 3\delta.$
It is thus a uniform $3\delta$-representation. Assume further that $\left\{\tau\left(k\right)\right\}_{k=1}^{n}$ generate a matrix $M$ with $rank\left(M-I_{n}\right)=n$, thus
$\varphi\left(abab^{2}\cdots ab^{t}\right)=\tau\left(1\right)\cdots\tau\left(t\right)=M,$
for some $t\in\mathbb{N}$.
Let $\psi:\Gamma\to GL_{N}\left(F\right)$ be a representation and $\varepsilon_{0}=d^{flex}\left(\varphi,\psi\right)$. Since $rk\left(\varphi\left(a\right)-I_{n}\right)=0$ and $rk\left(\varphi\left(b\right)-I_{n}\right)\leq\delta$, we have
$rk\left(\psi\left(a\right)-I_{N}\right)$ $\leq rk\left(\psi\left(a\right)-\varphi\left(a\right)\right)+rk\left(\varphi\left(a\right)-I_{n}\right)+rk\left(I_{n}-I_{N}\right)\leq 2\varepsilon_{0},$
$rk\left(\psi\left(b\right)-I_{N}\right)$ $\leq rk\left(\psi\left(b\right)-\varphi\left(b\right)\right)+rk\left(\varphi\left(b\right)-I_{n}\right)+rk\left(I_{n}-I_{N}\right)\leq 2\varepsilon_{0}+\delta.$
Since $\psi$ is a representation, any $v\in W\vcentcolon=ker\left(\psi\left(a\right)-I_{N}\right)\cap ker\left(\psi\left(b\right)-I_{N}\right)$ is fixed by any word in $\psi\left(a\right),\psi\left(b\right)$. We have
$dim\left(W\right)$ $\geq N-rank\left(\psi\left(a\right)-I_{N}\right)-rank\left(\psi\left(b\right)-I_{N}\right)$
$\geq N-\left(4\varepsilon_{0}+\delta\right)n.$
Thus for any $w\in\mathbb{F}_{2}$,
$rank\left(\psi\left(w\right)-I_{N}\right)$ $\leq N-dim\left(W\right)\leq\left(4\varepsilon_{0}+\delta\right)n$
$\implies rk\left(\psi\left(abab^{2}\cdots ab^{t}\right)-I_{N}\right)\leq 4\varepsilon_{0}+\delta.$
But by construction we also have
$rk\left(\varphi\left(abab^{2}\cdots ab^{t}\right)-I_{n}\right)=1$
$\implies rk\left(\psi\left(abab^{2}\cdots ab^{t}\right)-I_{N}\right)\geq$ $1-rk\left(\psi\left(abab^{2}\cdots ab^{t}\right)-\varphi\left(abab^{2}\cdots ab^{t}\right)\right)-rk\left(I_{N}-I_{n}\right)$
$\geq$ $1-2\varepsilon_{0},$
and hence
$1-2\varepsilon_{0}\leq 4\varepsilon_{0}+\delta$
$\implies\varepsilon_{0}\geq\frac{1}{6}-\frac{\delta}{6}=\frac{1}{6}-\frac{1}{6n}.$
proof of Theorem 1.3.
Let $\pi:\Gamma\to\mathbb{F}_{2}$ be a surjection, and choose $\gamma_{1},\gamma_{2}\in\Gamma$ such that $\pi\left(\gamma_{1}\right)=a$ and $\pi\left(\gamma_{2}\right)=b$. Define $\tilde{\varphi}\vcentcolon=\varphi\circ\pi:\Gamma\to GL_{n}\left(F\right)$ where $\varphi$ is as in Proposition 4.1. Then
$rk\left(\tilde{\varphi}\left(\gamma_{1}\right)-I_{n}\right)$ $=0$
$rk\left(\tilde{\varphi}\left(\gamma_{2}\right)-I_{n}\right)$ $\leq\frac{1}{n},$
and there is a word $w\in\mathbb{F}_{2}$ such that
$rk\left(\tilde{\varphi}\left(w\left(\gamma_{1},\gamma_{2}\right)-I_{n}\right)\right)=1,$
where $w\left(\gamma_{1},\gamma_{2}\right)$ is the word substitution in $\Gamma$. Arguing similarly to the proof of Proposition 4.1, we have for any representation $\psi:\Gamma\to GL_{N}\left(F\right)$,
$1-2d^{flex}\left(\tilde{\varphi},\psi\right)\leq rk\left(w\left(\gamma_{1},\gamma_{2}\right)-I_{N}\right)\leq 4d^{flex}\left(\tilde{\varphi},\psi\right)+\frac{1}{n}$
$\implies d^{flex}\left(\tilde{\varphi},\psi\right)\geq\frac{1}{6}-\frac{1}{6n}.$
Remark 4.2.
Choosing $\left\{\tau\left(k\right)\right\}_{k=1}^{n}$ to be transpositions $\left(k\quad k+1\right)$ in $S_{n}$ gives another example witnessing the instability of $\mathbb{F}_{2}$ in permutations, in addition to [4], that is not close to representation in any field $F$, and in particular is bounded from homomorphisms to $S_{n}$.